# Which Optimizer Works Best for Physics-Informed Neural Networks and Kolmogorov-Arnold Networks?

### v4

@@ Abstract
Physics-Informed Neural Networks (PINNs) have revolutionized the computation of PDE solutions by integrating partial differential equations (PDEs) into the neural network’s training process as soft constraints, becoming an important component of the scientific machine learning (SciML) ecosystem. More recently, physics-informed Kolmogorv-Arnold networks (PIKANs) have also shown to be effective and comparable in accuracy with PINNs. In their current implementation, both PINNs and PIKANs are mainly optimized using first-order methods like Adam, as well as quasi-Newton methods such as BFGS and its low-memory variant, L-BFGS. However, these optimizers often struggle with highly non-linear and non-convex loss landscapes, leading to challenges such as slow convergence, local minima entrapment, and (non)degenerate saddle points. In this study, we investigate the performance of Self-Scaled BFGS (SSBFGS), Self-Scaled Broyden (SSBroyden) methods and other advanced quasi-Newton schemes, including BFGS and L-BFGS with different line search strategies. These methods dynamically rescale updates based on historical gradient information, thus enhancing training efficiency and accuracy. We systematically compare these optimizers – using both PINNs and PIKANs – on key challenging PDEs, including the Burgers, Allen-Cahn, Kuramoto-Sivashinsky, Ginzburg-Landau, and Stokes equations. Additionally, we evaluate the performance of SSBFGS and SSBroyden for Deep Operator Network (DeepONet) architectures, demonstrating their effectiveness for data-driven operator learning. Our findings provide state-of-the-art results with orders-of-magnitude accuracy improvements without the use of adaptive weights or any other enhancements typically employed in PINNs. More broadly, our results reveal insights into the effectiveness of second-order optimization strategies in significantly improving the convergence and accurate generalization of PINNs and PIKANs.
@@ 1 Introduction
This section begins by providing the necessary background and motivation for our study, highlighting the key challenges and objectives. Following this, we present an overview of the BFGS (Broyden–Fletcher–Goldfarb–Shanno), SSBFGS (Self-Scaled BFGS), and SSBroyden (Self-Scaled Broyden) optimization methods, discussing their theoretical foundations and main characteristics. Section 2 presents a comprehensive study comparing different optimization methods with various line search and trust-region strategies. Specifically, in Subsections 2.1.1 and 2.1.2, we discuss a wide range of scenarios for the Burgers equation. We extend our study to more complex PDEs, including the Allen–Cahn equation in Section 2.3, the Kuramoto–Sivashinsky equation in Subsection 2.4, and the Ginzburg–Landau equation in Subsection 2.5. To demonstrate the efficacy of the proposed optimizers, we also solve the Stokes equation for viscous flow in a lid-driven wedge in Subsection 2.6. Furthermore, the performance of SSBFGS and SSBroyden is investigated for data-driven operator learning using Deep Operator Network (DeepONet) in Section 3. The conclusions of our study are summarized in Section 4. Finally, in Appendix A, we review the performance of the optimizers on the Lorenz system. In Appendix B, we provide a quantitative analysis of BFGS and SSBroyden on the multi-dimensional Rosenbrock function, serving as a pedagogical example to illustrate the effect of dimensionality on optimization performance.
@@ 1.1 Background and Motivation
Currently, the standard optimization algorithms used in PINNs are Adam and quasi-Newton methods, such as BFGS or its low-memory variant, L-BFGS. BFGS and L-BFGS enhance significantly the performance of PINNs by incorporating second-order information of the trainable variable space to precondition it. Although these two optimizers are very commonly employed not only in PINNs but also in other optimization problems because of their theoretical and numerical properties, recent experience in PINNs has shown that other algorithms could outperform them for a great variety of problems [51, 52, 53]. Among all of them we can highlight the Self-Scaled Broyden algorithms [54], which update the approximation to the inverse Hessian matrix using a self-scaling technique. This advanced optimization technique enhances convergence by adaptively scaling gradient updates based on historical error information. Furthermore, [55] investigates self-scaling within quasi-Newton methods from the Broyden family, introducing innovative scaling schemes and updates. Urban et al. [53] further demonstrated that these advanced optimizers, when combined with rescaled loss functions, significantly enhance the efficiency and accuracy of PINNs, enabling smaller networks to solve problems with significantly greater accuracy. As PINNs often suffer from conflicting gradients between multiple loss components (e.g., data and physics losses), resolving these conflicts is essential for stable and accurate training. Recent work by Wang et al. [56] demonstrates that second-order optimizers, particularly Newton-type methods like SOAP, exhibit positive gradient alignment and can effectively mitigate such conflicts, leading to improved convergence in multi-task optimization settings, like PINNs. Second-order optimization is also discussed in [57], where the authors present an implementation of a second-order optimizer and demonstrate improvements in both step time and wall-clock time across multiple large-scale tasks spanning diverse domains.
@@
Figure 1: Loss landscape of $\mathcal{L}=\mathcal{L}_{\text{Data}}+\lambda\mathcal{L}_{\text{PDE}}$ for the viscous Burgers equation (16). In the left subfigure, the landscape is shown for $\lambda=0$. This loss landscape is relatively smooth and exhibits well-defined convexity. However, it only fits the data (initial and boundary conditions) and does not converge to the solution within the domain. In contrast, the right subfigure illustrates the loss landscape for $\lambda=1.0$. Unlike the smooth landscape in the left subfigure, this landscape contains numerous local minima and saddle points, potentially leading to (non)-degenerate saddle points. The loss landscape is computed by using the method proposed by [60].
To provide an intuitive explanation, Figure 1 illustrates the loss landscape $\mathcal{L}=\mathcal{L}_{D}+\lambda\mathcal{L}_{PDE}$ for the viscous Burgers Equation (16), using the method proposed by [60]. The left subfigure of Figure 1 corresponds to $\lambda=0$, considering only the initial and boundary conditions. This loss landscape is smooth and exhibits well-defined convexity. However, it fits only the initial and boundary condition data and fails to converge to the solution within the domain. In contrast, the right subfigure depicts the loss landscape for $\lambda=1.0$. Unlike the smooth landscape on the left, this landscape is crinkled with numerous local minima and saddle points, including potentially degenerate and non-degenerate saddle points [61]. Consequently, the presence of multiple local minima can trap the optimizer, making it extremely challenging to minimize the loss function effectively. Motivated by this, in this work, we propose and analyze a suite of second-order optimizers and their applicability to linear, nonlinear, stiff, and chaotic PDEs whose solutions are computed by PINN and PIKAN types of approximation. Furthermore, we introduce a set of optimizers tailored to various classes of PDEs, achieving unprecedented convergence to machine precision with both single and double precision arithmetic. To the best of our knowledge, the current study is the most systematic and comprehensive investigation of how optimizers affect the performance of both PINNs and PIKANs.
@@ 1.2 Overview of BFGS and SSBroyden
In optimization, a fundamental challenge is determining the optimal direction and step size to transition from the current point $\mathbf{x}_{k}$ to an improved solution. In general, two primary approaches address this problem: line search methods and trust region methods. Both rely on a quadratic model to approximate the objective function around the current iterate. The key distinction lies in how they utilize this approximation; line-search methods determine the step size along a chosen direction, while trust-region methods restrict the step to a pre-defined neighborhood where the model is considered reliable.
1.2.1 Line-search methods
$\min_{\alpha>0}f(\mathbf{x}_{k}+\alpha\mathbf{p}_{k}).$ (1)
$\mathbf{x}_{k+1}=\mathbf{x}_{k}+\alpha_{k}\mathbf{p}_{k}.$ (2)
Wolfe conditions
which commonly receive the names of Armijo and sufficient decrease conditions, respectively. The quantities $c_{1}$ and $c_{2}$ are constants that should follow $10<c_{1}<c_{2}<$ 1. While the first condition ensures a sufficient decrease in the function $f$, the second one rules out unacceptably short steps. The latter inequality is commonly replaced by taking absolute values at both sides
Backtracking
$\textbf{p}_{k}^{\top}\nabla f_{k}\leq 0.$ (6)
1.2.2 Trust-region methods
Another class of optimization methods is the family of trust-region methods. The key distinction between trust-region and line-search methods is that the former defines a region (the trust region) around the current point $\mathbf{x}_{k}$ and seeks a suitable point within this region that reduces the objective function. To achieve this, a trust-region algorithm first constructs a local quadratic approximation of the function, denoted as $m_{k}$, and then approximately solves the following subproblem:
subject to:
$\|p\|\leq\Delta_{k},$
$\textbf{x}_{k+1}=\textbf{x}_{k}+\textbf{p}_{k},$
that is, in a trust-region algorithm the direction and the step-length are calculated at the same time.
@@ 1.2.3 Quasi-Newton methods
$f(x,y)=x^{2}+y^{2}+xy,$
in Figures 2 and 3. We see that from iteration 1 to 6, the algorithm iteratively predicts the search direction. In this example, the trust-region radius is kept constant since $f(x)$ is convex. However, for PDE problems, we have implemented an adaptive trust-region size.
The choice between backtracking line-search and the trust-region method depends on the problem characteristics and specific implementation details. Backtracking line-search is often preferable for simpler or computationally inexpensive problems due to its straightforward nature and adaptability. In contrast, trust-region methods are more suitable for complex or constrained optimization tasks where careful control over the step size is essential. A key distinction is that backtracking line-search directly adjusts the step size without solving quadratic subproblems, making it computationally faster in cases where objective function and gradient evaluations are inexpensive. Ultimately, the choice between these methods depends on the specific problem and computational trade-offs.
$\|\nabla f(\bm{x}_{k})\|\leq\epsilon,$ (9)
  1. 1.
  2. 2.
    If $\|\bm{x}_{k}-\bm{x}_{k-1}\|\leq\eta$, where $\eta$ is a small tolerance.
  3. 3.
@@ 1.3 Broyden Family of Optimizers
$\tau_{k}$ =$=\begin{cases}\min\left(\rho_{k}^{+}\sigma_{k}^{(1-N)},\sigma_{k}\right),&\text{if }\theta_{k}\leq 0,\\[6.0pt] \rho_{k}^{+}\min\left(\sigma_{k}^{(1-N)},\frac{1}{\theta_{k}}\right),&\text{otherwise}.\end{cases}$
@@
In the context of Physics-Informed Neural Networks (PINNs) based on Multi-Layer Perceptrons (MLPs), recent research has begun to explore various families of SSBroyden updates and modifications to the loss formulation to address known challenges such as gradient conflicts and ill-conditioning. These second-order strategies significantly enhance both convergence speed and prediction accuracy compared to classical BFGS. In this work, we extend these findings to more complex problems beyond those presented in [53], employing not only SSBroyden but also other techniques proposed in the PINN literature. Additionally, we investigate whether these BFGS modifications enhance the performance of the recently introduced Kolmogorov-Arnold Networks (KANs) [3]. This study will further enable fair comparisons between KANs and MLPs using more advanced optimization algorithms. In the following section, we present a comprehensive study comparing the performance of different optimization methods with various line-search and trust-region strategies for solving PDEs. For each example, we analyze and discuss the performance of both PINNs and PIKANs.
@@ 2.1 Burgers equation
and periodic boundary conditions. The viscosity is given by $\nu=\frac{0.01}{\pi}\approx 0.003$, and the spatio-temporoal domain for computing the solution is $(x,\leavevmode\nobreak\ t)\in[-1,1]\times[0,1]$.
@@ 2.1.1 BFGS and SSBroyden optimization with Strong Wolfe line-search
Figure 4: PINNs with single precision for the Burgers equation: The evolution of the loss function (left) and corresponding $l_{2}$ relative errors (right) over iterations for Case 1. After approximately 4,000 iterations, the loss function and relative error stabilize, showing no significant further improvement.
1 Adam [1000] + BFGS with Wolfe [10000] $1.04\times 10^{-4}$ 263 1,341
1 Adam [1000] + SSBroyden with Wolfe [10000] $3.60\times 10^{-5}$ 209 1,341
1 Adam [1000] + SSBFGS with Wolfe [10000] $4.04\times 10^{-5}$ 179 1,341
Table 1: PINNs with single precision for the Burgers equation: Relative L2 error and training time for solving the Burgers equation using single precision with different optimizers. A PINN with four layers, each containing 20 neurons, is trained for 1000 iterations using the Adam optimizer, followed by 10,000 iterations with the specified optimizer.
Figure 5: Double-Precision PINNs for the Burgers equation: The evolution of the loss function (left) and corresponding $l_{2}$ relative errors (right) over iterations for Case 2. After approximately 4,000 iterations, the loss function and relative error stabilize, showing no significant further improvement.
@@
Table 2: Double-Precision PINNs for the Burgers equation: Case 2: A PINN with four layers, each containing 20 neurons, is trained for 1000 iterations using the Adam optimizer, followed by 50,000 iterations with BFGS, SSBroyden, SSBFGS, and L-BFGS optimizers with Strong Wolfe line-search. Case 3: The same network structure as Case 2 is used, but the number of iterations with BFGS, and SSBroyden is reduced to 30,000. Case 4: The same network structure as in Case 2 and Case 3 is used; however, instead of starting with the Adam optimizer, the training begins directly with second-order optimizers, namely BFGS, SSBFGS, and SSBroyden. Case 5: A deeper PINN with eight layers (20 neurons per layer) is trained for 1000 iterations using Adam, followed by 30,000 iterations with BFGS and SSBroyden.
Figure 6: Double-Precision PIKANs for the Burgers Equation: Evolution of the loss function (left) and $l_{2}$ relative errors (right) for Case 6. The results compare the performance of PIKANs trained with SSBroyden and BFGS optimizers.
@@
Table 3: Double-Precision PIKANs for the Burgers Equation: Relative $l_{2}$ error and training time for solving the Burgers equation using double precision with different optimizers. Case 6 highlights a PIKAN with four layers of 20 neurons each and degree 3. A PIKAN with three layers, each containing 10 neurons, is trained using Chebyshev polynomials of degree 3 in Case 7 and degree 5 in Case 8. The training consists of 1000 iterations with the Adam optimizer, followed by 50,000 iterations with BFGS and SSBroyden optimizers. Case 6 demonstrates improved performance despite a significant increase in training time.
Figures 4 illustrate the loss function and relative error over iterations for Case 1 in single precision. The results are obtained using a PINN with four layers, each comprising 20 neurons, trained with the Adam optimizer followed by 10,000 iterations of the specified optimizer. Table 1 provides a summary of the corresponding relative error and training time. Notably, SSBFGS and SSBroyden achieve a relative error of $10^{-5}$, while BFGS converges to approximately $10^{-4}$. It is observed that the loss function and relative error stabilize after roughly 4,000 iterations, indicating no significant further improvement.
@@
The performance of PINNs with a similar network architecture and double precision is presented in Table 2 for Case 2 and Case 3. The same network architecture is employed, consisting of four hidden layers, each with 20 neurons. After the initial Adam optimization stage, the training proceeds with BFGS, SSBFGS, and SSBroyden optimizers for two different iteration counts: Case 2 utilizes 50,000 iterations, while Case 3 uses 30,000 iterations. In Case 4, the same network structure as in Case 2 and Case 3 is used; however, training begins directly with second-order optimizers, BFGS, SSBFGS, and SSBroyden, instead of the Adam optimizer. The results, as shown in the table, demonstrate that increasing the iteration count to 50,000 significantly raises the training time. While this reduces SSBFGS and SSBroyden’s error to $10^{-8}$, it does not result in any noticeable improvement in the error achieved by BFGS. In Case 5, a deeper network with eight hidden layers is used, trained for 30,000 iterations. This configuration further improves the error, though at the cost of a substantial increase in training time. Figure 5 illustrates the loss function and relative error over iterations for Case 2 using four optimizers, BFGS, SSBFGS, SSBroyden, and L-BFGS. As summarized in Table 2, the relative errors achieved are $10^{-8}$ for SSBroyden, $10^{-5}$ for BFGS, and $10^{-3}$ for L-BFGS.
@@
Additionally, Figure 6 compares the performance of BFGS and SSBroyden for PIKANs in Case 6. A PIKAN with four layers, each containing 20 neurons, is trained using Chebyshev polynomials of degree 3. As reported in Table 3, the error in this case is very close to that of Case 2 in Table 2. Furthermore, a PIKAN with three layers, each containing 10 neurons, is trained using Chebyshev polynomials of degree 3 in Case 7 and degree 5 in Case 8. The results demonstrate that SSBroyden consistently outperforms BFGS. Although the networks are trained for 50,000 iterations, minimal improvement is observed beyond approximately 30,000 iterations. As summarized in Table 3, increasing the polynomial degree from 3 to 5 does not improve the results, while the training time nearly doubles with the higher degree.
@@
The conclusion highlights that the best error for the Burgers equation was achieved using SSBroyden with double precision. Across all seven case studies, SSBroyden consistently outperformed both BFGS and L-BFGS. A comparison between PINNs and PIKANs reveals that, for the same number of parameters, PIKANs did not perform as well. However, when trained with a higher number of parameters—at the cost of significantly increased training time—PIKANs can achieve results comparable to PINNs. In general, a comparison between Case 2 and Case 6 shows that, although both achieve nearly identical errors, PINNs with SSBroyden demonstrate more efficient training times compared to PIKANs.

### v5

# Optimizing the Optimizer for Physics-Informed Neural Networks and Kolmogorov-Arnold Networks

@@ Abstract
Physics-Informed Neural Networks (PINNs) have revolutionized the computation of PDE solutions by integrating partial differential equations (PDEs) into the neural network’s training process as soft constraints, becoming an important component of the scientific machine learning (SciML) ecosystem. More recently, physics-informed Kolmogorv-Arnold networks (PIKANs) have also shown to be effective and comparable in accuracy with PINNs. In their current implementation, both PINNs and PIKANs are mainly optimized using first-order methods like Adam, as well as quasi-Newton methods such as BFGS and its low-memory variant, L-BFGS. However, these optimizers often struggle with highly non-linear and non-convex loss landscapes, leading to challenges such as slow convergence, local minima entrapment, and (non)degenerate saddle points. In this study, we investigate the performance of Self-Scaled BFGS (SSBFGS), Self-Scaled Broyden (SSBroyden) methods and other advanced quasi-Newton schemes, including BFGS and L-BFGS with different line search strategies. These methods dynamically rescale updates based on historical gradient information, thus enhancing training efficiency and accuracy. We systematically compare these optimizers – using both PINNs and PIKANs – on key challenging PDEs, including the Burgers, Allen-Cahn, Kuramoto-Sivashinsky, Ginzburg-Landau, and Stokes equations. Additionally, we evaluate the performance of SSBFGS and SSBroyden for Deep Operator Network (DeepONet) architectures, demonstrating their effectiveness for data-driven operator learning. Our findings provide state-of-the-art results with orders-of-magnitude accuracy improvements without the use of adaptive weights or any other enhancements typically employed in PINNs. More broadly, our results reveal insights into the effectiveness of quasi-Newton optimization strategies in significantly improving the convergence and accurate generalization of PINNs and PIKANs.
@@ 1 Introduction
This section begins by providing the necessary background and motivation for our study, highlighting the key challenges and objectives. Following this, we present an overview of the BFGS (Broyden–Fletcher–Goldfarb–Shanno), SSBFGS (Self-Scaled BFGS), and SSBroyden (Self-Scaled Broyden) optimization methods, discussing their theoretical foundations and main characteristics. Section 2 presents a comprehensive study comparing different optimization methods with various line search and trust-region strategies. Specifically, in Subsections 2.1.1 and 2.1.2, we discuss a wide range of scenarios for the Burgers equation. We extend our study to more complex PDEs, including the Allen–Cahn equation in Section 2.3, the Kuramoto–Sivashinsky equation in Subsection 2.4, and the Ginzburg–Landau equation in Subsection 2.5. To demonstrate the efficacy of the proposed optimizers, we also solve the Stokes equation for viscous flow in a lid-driven wedge in Subsection 2.6. Furthermore, the performance of SSBFGS and SSBroyden is investigated for data-driven operator learning using Deep Operator Network (DeepONet) in Section 3. The conclusions of our study are summarized in Section 4. Finally, in Appendix C, we review the performance of the optimizers on the Lorenz system. In Appendix D, we provide a quantitative analysis of BFGS and SSBroyden on the multi-dimensional Rosenbrock function, serving as a pedagogical example to illustrate the effect of dimensionality on optimization performance.
@@ 1.1 Background and Motivation
Currently, the standard optimization algorithms used in PINNs are Adam and quasi-Newton methods, such as BFGS or its low-memory variant, L-BFGS. BFGS and L-BFGS enhance significantly the performance of PINNs by incorporating second-order information of the trainable variable space to precondition it. Although these two optimizers are very commonly employed not only in PINNs but also in other optimization problems because of their theoretical and numerical properties, recent experience in PINNs has shown that other algorithms could outperform them for a great variety of problems [51, 52, 53]. Among all of them we can highlight the Self-Scaled Broyden algorithms [54], which update the approximation to the inverse Hessian matrix using a self-scaling technique. This advanced optimization technique enhances convergence by adaptively scaling gradient updates based on historical error information. Furthermore, [55] investigates self-scaling within quasi-Newton methods from the Broyden family, introducing innovative scaling schemes and updates. Urban et al. [53] further demonstrated that these advanced optimizers, when combined with rescaled loss functions, significantly enhance the efficiency and accuracy of PINNs, enabling smaller networks to solve problems with significantly greater accuracy. As PINNs often suffer from conflicting gradients between multiple loss components (e.g., data and physics losses), resolving these conflicts is essential for stable and accurate training. Recent work by Wang et al. [56] demonstrates that quasi-Newton optimizers, particularly Newton-type methods like SOAP, exhibit positive gradient alignment and can effectively mitigate such conflicts, leading to improved convergence in multi-task optimization settings, like PINNs. Quasi-Newton optimization is also discussed in [57], where the authors present an implementation of a quasi-Newton optimizer and demonstrate improvements in both step time and wall-clock time across multiple large-scale tasks spanning diverse domains.
@@
Figure 1: Projection of the loss landscape $\mathcal{L}=\mathcal{L}_{\text{Data}}+\lambda\mathcal{L}_{\text{PDE}}$ for the viscous Burgers equation (9) onto a 2D parameter subspace extracted from SVD of global parameter space, as described in [60]. The vertical axis represents the scalar loss value $\mathcal{L}(\theta)$, while the horizontal axes correspond to randomly chosen linear combinations of network parameters. In the left subfigure, the landscape is shown for $\lambda=0$, where only data loss is used. The resulting landscape is relatively smooth and convex but fails to enforce the PDE constraints within the domain. In contrast, the right subfigure shows the landscape for $\lambda=1.0$, where the PDE residual is included. This introduces a highly non-convex landscape with numerous local minima and saddle points, reflecting the increased stiffness and complexity introduced by the physics term.
To provide an intuitive explanation, Figure 1 illustrates the loss landscape $\mathcal{L}=\mathcal{L}_{D}+\lambda\mathcal{L}_{PDE}$ for the viscous Burgers Equation (9), using the method proposed by [60]. The left subfigure of Figure 1 corresponds to $\lambda=0$, considering only the initial and boundary conditions. This loss landscape is smooth and exhibits well-defined convexity. However, it fits only the initial and boundary condition data and fails to converge to the solution within the domain. In contrast, the right subfigure depicts the loss landscape for $\lambda=1.0$. Unlike the smooth landscape on the left, this landscape is crinkled with numerous local minima and saddle points, including potentially degenerate and non-degenerate saddle points [61]. Consequently, the presence of multiple local minima can trap the optimizer, making it extremely challenging to minimize the loss function effectively. Motivated by this, in this work, we propose and analyze a suite of quasi-Newton optimizers and their applicability to linear, nonlinear, stiff, and chaotic PDEs whose solutions are computed by PINN and PIKAN types of approximation. Furthermore, we introduce a set of optimizers tailored to various classes of PDEs, achieving unprecedented convergence to machine precision with both single and double precision arithmetic. To the best of our knowledge, the current study is the most systematic and comprehensive investigation of how optimizers affect the performance of both PINNs and PIKANs.
@@ 1.2 Overview of BFGS and SSBroyden
In optimization, a fundamental challenge is determining the optimal direction and step size to transition from the current point $\mathbf{x}_{k}$ to an improved solution. In general, two primary approaches address this problem: line-search methods and trust-region methods. Both rely on a quadratic model to approximate the objective function around the current iterate. The key distinction lies in how they utilize this approximation; line-search methods determine the step size along a chosen direction, while trust-region methods restrict the step to a pre-defined neighborhood where the model is considered reliable. The details of the line-search and trust-region methods are provided in A.
@@ 1.3 Broyden Family of Optimizers
$\tau_{k}$ $=\begin{cases}\min\left(\rho_{k}^{+}\sigma_{k}^{(1-N)},\sigma_{k}\right),&\text{if }\theta_{k}\leq 0,\\[6.0pt] \rho_{k}^{+}\min\left(\sigma_{k}^{(1-N)},\frac{1}{\theta_{k}}\right),&\text{otherwise}.\end{cases}$
@@
In the context of Physics-Informed Neural Networks (PINNs) based on Multi-Layer Perceptrons (MLPs), recent research has begun to explore various families of SSBroyden updates and modifications to the loss formulation to address known challenges such as gradient conflicts and ill-conditioning. These second-order strategies significantly enhance both convergence speed and prediction accuracy compared to classical BFGS. In this work, we extend these findings to more complex problems beyond those presented in [53], employing not only SSBroyden but also other techniques proposed in the PINN literature. Additionally, we investigate whether these BFGS modifications enhance the performance of the recently introduced Kolmogorov-Arnold Networks (KANs) [3]. This study will further enable fair comparisons between KANs and MLPs using more advanced optimization algorithms.
In addition to PINNs, we also evaluate the performance of quasi-Newton optimization methods in the context of Deep Operator Networks (DeepONets) [75]. DeepONet is a neural architecture designed to learn nonlinear operators—mappings between infinite-dimensional function spaces—and has shown strong performance in solving parametric PDEs in purely data-driven settings. While PINNs incorporate the governing physics directly into the loss function, DeepONets rely solely on data to learn the underlying operator. We demonstrate that the benefits of quasi-Newton methods are not limited to physics-informed frameworks, but also extend to operator learning tasks. This inclusion broadens the impact of our work by showing that improved optimization strategies can enhance training stability and accuracy across diverse neural PDE solvers. This extension allows us to evaluate whether quasi-Newton optimizers like SSBroyden can consistently improve training performance across different PDE learning paradigms.
In the following section, we present a comprehensive study comparing the performance of different optimization methods with various line-search and trust-region strategies for solving PDEs. For each example, we analyze and discuss the performance of both PINNs and PIKANs.
@@ 2.1 Burgers equation
and periodic boundary conditions. The viscosity is given by $\nu=\frac{0.01}{\pi}\approx 0.003$, and the spatio-temporoal domain for computing the solution is $(x,\leavevmode\nobreak\ t)\in[-1,1]\times[0,1]$. In the following subsections, we present the performance of L-BFGS, BFGS, SSBFGS, and SSBroyden optimizers using Strong Wolfe line-search, backtracking, and trust-region strategies (see Subsections 2.1.1 and 2.1.2). All results in these sections are reported in double precision. To provide a more comprehensive study, we also include single-precision results in B.
@@ 2.1.1 BFGS and SSBroyden optimization with Strong Wolfe line-search
Figure 2: Double-Precision PINNs for the Burgers equation: The evolution of the loss function (left) and corresponding $l_{2}$ relative errors (right) over iterations for Case 1.
@@
Table 1: Double-Precision PINNs for the Burgers equation: Case 1: A PINN with four layers, each containing 20 neurons, is trained for 1000 iterations using the Adam optimizer, followed by 50,000 iterations with BFGS, SSBroyden, SSBFGS, and L-BFGS optimizers with Strong Wolfe line-search. Case 2: The same network structure as Case 1 is used, but the number of iterations with BFGS, and SSBroyden is reduced to 30,000. Case 3: The same network structure as in Case 1 and Case 2 is used; however, instead of starting with the Adam optimizer, the training begins directly with quasi-Newtonoptimizers, namely BFGS, SSBFGS, and SSBroyden. Case 4: A deeper PINN with eight layers (20 neurons per layer) is trained for 1000 iterations using Adam, followed by 30,000 iterations with BFGS and SSBroyden.
Figure 3: Double-Precision PIKANs for the Burgers Equation: Evolution of the loss function (left) and $l_{2}$ relative errors (right) for Case 5. The results compare the performance of PIKANs trained with SSBroyden and BFGS optimizers.
@@
Table 2: Double-Precision PIKANs for the Burgers Equation: Relative $l_{2}$ error and training time for solving the Burgers equation using double precision with different optimizers. Case 5 highlights a PIKAN with four layers of 20 neurons each and degree 3. A PIKAN with three layers, each containing 10 neurons, is trained using Chebyshev polynomials of degree 3 in Case 6 and degree 5 in Case 7. The training consists of 1000 iterations with the Adam optimizer, followed by 50,000 iterations with BFGS and SSBroyden optimizers. Case 5 demonstrates improved performance despite a significant increase in training time.
@@
The performance of PINNs for Case 1 and Case 2 is presented in Table 1. Both cases use the same network architecture, consisting of four hidden layers with 20 neurons each, and are evaluated in double precision. After the initial Adam optimization stage, the training proceeds with BFGS, SSBFGS, and SSBroyden optimizers for two different iteration counts: Case 1 utilizes 50,000 iterations, while Case 2 uses 30,000 iterations. In Case 3, the same network structure as in Case 1 and Case 2 is used; however, training begins directly with quasi-Newton optimizers, BFGS, SSBFGS, and SSBroyden, instead of the Adam optimizer. The results, as shown in the table, demonstrate that increasing the iteration count to 50,000 significantly raises the training time. While this reduces SSBFGS and SSBroyden’s error to $10^{-8}$, it does not result in any noticeable improvement in the error achieved by BFGS. In Case 4, a deeper network with eight hidden layers is used, trained for 30,000 iterations. This configuration further improves the error, though at the cost of a substantial increase in training time. Figure 2 illustrates the loss function and relative error over iterations for Case 1 using four optimizers, BFGS, SSBFGS, SSBroyden, and L-BFGS. As summarized in Table 1, the relative errors achieved are $10^{-8}$ for SSBroyden, $10^{-5}$ for BFGS, and $10^{-3}$ for L-BFGS.
@@
Additionally, Figure 3 compares the performance of BFGS and SSBroyden for PIKANs in Case 5. A PIKAN with four layers, each containing 20 neurons, is trained using Chebyshev polynomials of degree 3. As reported in Table 2, the error in this case is very close to that of Case 1 in Table 1. Furthermore, a PIKAN with three layers, each containing 10 neurons, is trained using Chebyshev polynomials of degree 3 in Case 6 and degree 5 in Case 7. The results demonstrate that SSBroyden consistently outperforms BFGS. Although the networks are trained for 50,000 iterations, minimal improvement is observed beyond approximately 30,000 iterations. As summarized in Table 2, increasing the polynomial degree from 3 to 5 does not improve the results, while the training time nearly doubles with the higher degree.
@@
The conclusion highlights that the best error for the Burgers equation was achieved using SSBroyden with double precision. Across all seven case studies, SSBroyden consistently outperformed both BFGS and L-BFGS. A comparison between PINNs and PIKANs reveals that, for the same number of parameters, PIKANs did not perform as well. However, when trained with a higher number of parameters—at the cost of significantly increased training time—PIKANs can achieve results comparable to PINNs. In general, a comparison between Case 1 and Case 5 shows that, although both achieve nearly identical errors, PINNs with SSBroyden demonstrate more efficient training times compared to PIKANs.
The performance of SSBroyden, SSBFGS, and BFGS using single-precision is presented in  B. It is important to note that the previous tables compare the performance of different optimizers under a fixed iteration budget. To further assess the efficiency of our proposed quasi-Newton methods, Table 3 presents results based on a fixed relative $L_{2}$ error threshold of $10^{-4}$. This setup enables a fair comparison of SSBFGS, SSBroyden, and BFGS in terms of both convergence speed and overall computational cost, independent of a predefined number of iterations. As shown in the table, both SSBFGS and SSBroyden reach the target accuracy significantly faster than standard BFGS, requiring fewer iterations and substantially less training time. Notably, SSBroyden achieves the target relative $L_{2}$ errors of $10^{-4}$ and $10^{-5}$ in just 1,307 and 2,864 iterations, respectively, compared to 5,621 and 69,332 iterations for BFGS. This results in a more than threefold reduction in training time. BFGS exhibits slow convergence and eventual stagnation in the later stages, suggesting that it struggles to make further progress. In contrast, the self-scaled variants (SSBFGS and SSBroyden) achieve the target error more efficiently and consistently.
Optimizer Iterations Final Rel. $L_{2}$ Error Training Time (s)
SSBFGS with Wolfe 1,567 $9.91\times 10^{-5}$ 167
SSBroyden with Wolfe 1,307 $8.98\times 10^{-5}$ 171
BFGS with Wolfe 5,621 $9.79\times 10^{-5}$ 586
SSBFGS with Wolfe 4,766 $9.97\times 10^{-6}$ 496
SSBroyden with Wolfe 2,864 $9.74\times 10^{-6}$ 310
BFGS with Wolfe 69,332 $9.99\times 10^{-6}$ 7189
Table 3: Comparison of optimizers for the Burgers equation with a target relative $L_{2}$ error of $10^{-4}$ and $10^{-5}$. A PINN with four layers, each containing 20 neurons, is trained directly using a quasi-Newton optimizer. Each optimizer is run until the error threshold is reached. The table reports the number of iterations, final relative error, and total training time. Both SSBFGS and SSBroyden achieve the target accuracy significantly faster than standard BFGS.